OpenStudy (anonymous):

Solve for each system of equations 2x + 3y = 10, 3x + 2y = 5

OpenStudy (anonymous):

u know how to solve it?

OpenStudy (anonymous):

No idea.

OpenStudy (anonymous):

there are two way to do it, subtitution and elimination u know any of it?

OpenStudy (anonymous):

nope, i never learned any of it. I was homeschooled from 5th-12th grade. and it was some stupid books that taught me nothing so i've been tying to teach myself everything on my own, and yeah.

OpenStudy (anonymous):

okay, so substitution is just changing one of the variable value, like your previous question

OpenStudy (anonymous):

and elimination is yeah it eliminate one of the variable

OpenStudy (anonymous):

\[2x+3y=10\]\[2x=-3y+10\]\[x=-\frac{ 3 }{ 2 }y +\frac{ 10 }{ 2 }\]

OpenStudy (anonymous):

that the substitute value of X, so just change X with that

OpenStudy (triciaal):

@ArifL 2 other methods are using matrix and graphing

OpenStudy (anonymous):

yeah, but it will be abit complex for matrix

OpenStudy (anonymous):

(4x - 6)(6x - 4) A. 24x^2 + 24x + 24 B. 24x^2 - 52x + 24 C. 24x^2 + 52x -24 D. None of these so what would this be?

OpenStudy (anonymous):

\[(ab) \times (cd) = ac+ad+bc+bd\]

OpenStudy (anonymous):

and thanks for reminding me @triciaal

OpenStudy (anonymous):

(-1,4)

OpenStudy (triciaal):

welcome

OpenStudy (anonymous):

and its B